### v1

# Efficient Reasoning for Large Reasoning Language Models via Certainty-Guided Reflection Suppression

Abstract
Recent Large Reasoning Language Models (LRLMs) employ long chain-of-thought reasoning with complex reflection behaviors, typically signaled by specific trigger words (e.g., “Wait” and “Alternatively”) to enhance performance. However, these reflection behaviors can lead to the overthinking problem where the generation of redundant reasoning steps that unnecessarily increase token usage, raise inference costs, and reduce practical utility. In this paper, we propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while maintaining reasoning accuracy. CGRS operates by dynamically suppressing the model’s generation of reflection triggers when it exhibits high confidence in its current response, thereby preventing redundant reflection cycles without compromising output quality. Our approach is model-agnostic, requires no retraining or architectural modifications, and can be integrated seamlessly with existing autoregressive generation pipelines. Extensive experiments across four reasoning benchmarks (i.e., AIME24, AMC23, MATH500, and GPQA-D) demonstrate CGRS’s effectiveness: it reduces token usage by an average of $18.5\%$ to $41.9\%$ while preserving accuracy. It also achieves the optimal balance between length reduction and performance compared to state-of-the-art baselines. These results hold consistently across model architectures (e.g., DeepSeek-R1-Distill series, QwQ-32B, and Qwen3 family) and scales (4B to 32B parameters), highlighting CGRS’s practical value for efficient reasoning.
Figure 1: Accuracy and output length of the Vanilla and proposed CGRS methods across different models, averaged over three mathematical reasoning benchmarks (AIME24, AMC23, and MATH500) and one scientific reasoning benchmark (GPQA-D). CGRS achieves an average token reduction of $18.5-41.9\%$ while maintaining performance.
1 Introduction
Large Reasoning Language Models (LRLMs), including OpenAI’s o1/o3 (OpenAI 2024, 2025) and DeepSeek-R1 (Guo et al. 2025), have demonstrated remarkable performance on demanding benchmarks, especially in advanced mathematics and program synthesis. A key factor behind their success is their slow-thinking capability, which combines step-by-step deduction with complex reflection behaviors, including backtracking, exploring alternative strategies, and self-verification of results  (Zhu and Li 2025).
Although these reflection behaviors enable models to self-correct and obtain accurate answers, they often suffer from a significant overthinking problem (Cuadron et al. 2025; Sui et al. 2025; Fan et al. 2025), where LRLMs persist in reasoning even after arriving at correct solutions. This phenomenon leads to substantial increases in token consumption, higher inference costs, and a degraded user experience due to unnecessary delays (Sun et al. 2025; Wang et al. 2025b). In extreme cases, excessively long responses may exceed context window limits, truncating critical information and compromising accuracy (Li et al. 2025a). Addressing this overthinking problem and developing efficient reasoning mechanisms, therefore represents a crucial challenge for improving LRLM performance and practicality.
To address this challenge, training-free methods have recently emerged as lightweight solutions that mitigate overthinking without requiring additional training (Zhu and Li 2025). For instance, prompt-guided approaches like TALE (Han et al. 2025) prompts model to respond within predefined token budgets through carefully designed instructions. While straightforward, these methods rely heavily on the model’s innate instruction-following capabilities. Decoding-manipulation techniques, such as Dynasor (Fu et al. 2024) and DEER (Yang et al. 2025b) dynamically adjust the decoding process to eliminate redundancy by terminating generation when the model reaches sufficient confidence. However, their effectiveness remains highly sensitive to the design of early-exit conditions. Consequently, current methods lack adaptive mechanisms to properly balance reflection suppression with reasoning quality, often either excessively suppressing valid corrections or inadequately curbing redundant reflections.
We propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while preserving reasoning accuracy. Previous work has demonstrated that reflection behaviors are typically signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (Muennighoff et al. 2025; Zhang et al. 2025; Ding et al. 2025). Our CGRS approach proactively suppresses the model’s generation of these reflection triggers when it exhibits high certainty in its current response, thereby preventing redundant reflection cycles, effectively alleviating overthinking, and preserving reasoning quality.
[…]
We evaluate CGRS across diverse open-source LRLMs spanning multiple architectures and scales: the DeepSeek-R1-Distill series (Qwen-7B and Llama-8B) (Guo et al. 2025), QwQ-32B (Team 2025), and the Qwen3 family (4B, 8B, 14B, and 32B) (Yang et al. 2025a). Our evaluation covers four reasoning benchmarks, including mathematical tasks (AIME24 (MAA Committees 2025), AMC23 (AI-MO 2024), and MATH500 (Lightman et al. 2023)) and scientific reasoning (GPQA Diamond (Rein et al. 2024)). As shown in Fig. 1, CGRS achieves significant token reduction ranging from $18.5\%$ to $41.9\%$, while maintaining comparable accuracy. Furthermore, CGRS outperforms existing efficient reasoning approaches, including both prompt-guided and decoding-manipulation methods, by delivering greater token efficiency with minimal accuracy impact. These results solidify CGRS as a superior solution for enabling efficient reasoning in LRLMs.
Our contributions are summarized as follows:
  • We propose CGRS, a training-free efficient reasoning method that alleviates the overthinking problem in LRLMs by dynamically suppressing reflection behaviors based on the model’s internal certainty signals;
  • We introduce two key components: (i) a certainty estimation mechanism that quantifies model confidence through entropy analysis of tentative answers, and (ii) a dynamic reflection trigger suppression technique that selectively disables unproductive reflection cycles;
  • We validate our method against state-of-the-art efficient reasoning methods spanning multiple model scales (4B to 32B parameters) on four reasoning benchmarks. Across all evaluations, CGRS demonstrates superior token efficiency with minimal impact on accuracy, positioning it as a superior solution for efficient reasoning in LRLMs.
2 Related Work
LRLMs have significantly advanced long Chain-of-Thought (CoT) reasoning through supervised fine-tuning and reinforcement learning (Pan et al. 2025). However, prolonged CoT often introduces redundant steps that increase computational costs, a phenomenon known as overthinking (Cuadron et al. 2025; Sui et al. 2025; Fan et al. 2025). To address this issue, many works focus on compressing verbose CoTs into more concise reasoning traces (Sui et al. 2025; Qu et al. 2025; Feng et al. 2025; Wang et al. 2025b; Zhu and Li 2025), which can be broadly categorized into two approaches: training-based methods (Arora and Zanette 2025; Aggarwal and Welleck 2025; Wang et al. 2025a; Ma et al. 2025b; Munkhbat et al. 2025; Chen et al. 2025b; Zhu et al. 2025) and training-free methods (Ma et al. 2025a; Yu, Yu, and Wang 2025; Han et al. 2025; Yang et al. 2025c; Team et al. 2025; Wu et al. 2025; Fu et al. 2024; Jiang et al. 2025; Yang et al. 2025b; Li et al. 2025b; Aytes, Baek, and Hwang 2025; Yang et al. 2025c). While effective, training-based methods often require retraining, careful data curation, and may compromise the general capabilities of the pretrained model. In this work, we focus on training-free methods, which eliminate the need for additional training and can be further divided into four categories: prompt-guided, pipeline-based, model-merging, and decoding-manipulation methods.
Prompt-guided Methods.
This type of method (Ma et al. 2025a; Yu, Yu, and Wang 2025; Han et al. 2025) compress lengthy CoT reasoning through prompt engineering. For instance, TALE (Han et al. 2025) simply prompts the model to solve the problem within token budgets. While simple, these methods heavily depend on the model’s inherent instruction-following capability.
Pipeline-based Methods.
This category of methods (Ong et al. 2025; Aytes, Baek, and Hwang 2025; Yang et al. 2025c) distribute queries or reasoning stages across multiple LRLMs, leveraging the efficiency of small models and the advanced reasoning of larger ones. However, these approaches introduce deployment complexity and additional inference overhead from auxiliary modules.
Model Merging.
Such methods (Team et al. 2025; Wu et al. 2025) combine slow- and fast-thinking models via weight averaging to enable long-to-short reasoning. While effective for medium-sized models, this approach underperforms at extreme scales (very large or small models) and lacks precise control over reasoning depth (Zhu and Li 2025).
Decoding-manipulation Methods.
This type of method (Fu et al. 2024; Jiang et al. 2025; Yang et al. 2025b; Li et al. 2025b) dynamically adjusts the decoding process to eliminate redundancy. For example, Dynasor (Fu et al. 2024) and DEER (Yang et al. 2025b) terminate decoding early when the model reaches sufficient confidence, though their performance depends critically on the exit condition design.
In this work, we focus on decoding manipulation for efficient reasoning. Our proposed CGRS method dynamically intervenes in token logits to suppress reflective behaviors based on the model’s internal certainty signals.
3 Method
In this section, we present our Certainty-Guided Reflection Suppression (CGRS) method for efficient reasoning in LRLMs.
3.1 Preliminaries
Large reasoning language models demonstrate unique generation patterns during inference. Given a problem $\mathbf{q}$} (with prompt included), the output $\mathbf{o}$} is generated token by token as ${\mathbf{o}}_{t}\sim{\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t})$, where $\bm{\pi}$} is the model. LRLMs organize their output using thinking delimiters (i.e., <think> and </think>), splitting the response into two key parts: slow thinking and conclusion. During slow thinking, LRLMs perform detailed, step-by-step reasoning before summarizing their thought process and delivering the final answer in the conclusion (Zhang et al. 2025).
Within slow thinking, models engage in complex reflection behaviors, including backtracking, exploring alternative approaches, and verifying results. These behaviors are often signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (Muennighoff et al. 2025; Zhang et al. 2025; Ding et al. 2025).
Notably, these reflection behaviors cause models to continue reasoning even after reaching an initial answer. They may perform extra validation steps or switch reasoning strategies in later stages, either to double-check previous results or explore different paths. However, this can sometimes lead to unproductive overthinking, where the model enters unnecessary reasoning loops, repeatedly validating or re-evaluating without meaningful progress (Chen et al. 2025a).
To address this, our CGRS method proactively suppresses the model’s tendency to generate reflection triggers when it exhibits high certainty in its current response. This prevents redundant reflection behavior and avoids overthinking. CGRS operates in two phases: (i) certainty estimation through checkpoint probing, and (ii) dynamic trigger suppression. We detail these components as follows.
3.2 Certainty Estimation
To estimate the necessity of reflective behaviors during inference, we introduce a certainty score that quantifies the LLM’s confidence in its currently generated response.
[…]
$C=1-\left(\frac{\frac{1}{n}\sum_{i=1}^{n}\mathcal{H}({\mathbf{p}}_{{\mathbf{a}}_{i}})}{\log(|{\mathbf{V}}|)}\right),$ (1)
where $n=|{\mathbf{a}}|$ is the token length of the probed result, ${\mathbf{p}}_{{\mathbf{a}}_{i}}={\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t},{\mathbf{c}},{\mathbf{a}}_{<i})$ represents the probability distribution of ${\mathbf{a}}_{i}$, $\mathcal{H}({\mathbf{p}})=-{\mathbf{p}}^{\top}\log({\mathbf{p}})$ denotes the token entropy, and $|{\mathbf{V}}|$ is the total number of tokens in the LLM’s vocabulary. The term $\log(|{\mathbf{V}}|)$ represents the maximum possible entropy (i.e., when the probability distribution is uniform across all tokens), serving as a normalization factor. This normalization bounds the certainty score $C$ to $[0,1]$, where a higher $C$ indicates higher LLM confidence (i.e., lower entropy in $\mathbf{a}$}), suggesting a reduced need for subsequent reflection behaviors.
Algorithm 1 The token prediction process in CGRS.
[…]
1: compute probability distribution ${\mathbf{p}}_{t}={\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t})$;
2: sample $r\sim\mathrm{Bernoulli}(p)$;
3:if $r=1$ then
4:  modify ${\mathbf{p}}_{t}$: set logits of tokens in $S_{\text{trigger}}$ to a large negative value and re-normalize;
5:end if
6: sample the next token ${\mathbf{o}}_{t}$ from ${\mathbf{p}}_{t}$;
7:if checkpoint marker appears then
8:  obtain tentative answers $\mathbf{a}$} given the input $({\mathbf{q}},{\mathbf{o}}_{<t},{\mathbf{c}})$;
9:  compute certainty score $C$ using $\mathbf{a}$} via Eq. (1);
10:  compute suppression probability $p$ using $C$ and $\delta$ via Eq. (2);
11:end if
12:return: the next token ${\mathbf{o}}_{t}$, the suppression probability pp.
3.3 Reflection Suppression
After estimating the LLM’s certainty regarding its current generation, the certainty score $C$ guides the suppression of subsequent reflection behaviors.
Reflection Triggers.
Following prior work (Yang et al. 2025c; Fu et al. 2024; Zhang et al. 2025), we consider three trigger categories: (i) core hesitation and transition words (e.g., “But” and “Wait”), (ii) alternative proposal markers (e.g., “Alternatively”), and (iii) colloquial contemplation cues (e.g., “Hmm”), all of which signal potential shifts in the LLM’s reasoning mode. Since a single trigger word has multiple natural language variants (e.g., “wait” as a lowercase variant of “Wait”) that map to different tokens in the vocabulary, we construct a comprehensive trigger token set $S_{\text{trigger}}$ through frequency analysis. Specifically, we identify all possible variants of each reflection trigger in the tokenizer’s vocabulary and then analyze their generation frequencies in reasoning traces. This analysis was performed using reasoning traces from the DeepSeek-R1-Distill-Qwen-7B model (Guo et al. 2025) on both the AIME24 (MAA Committees 2025) and AMC23 (AI-MO 2024) benchmarks. To ensure robustness, we conducted four independent inference runs per dataset, aggregating statistics across all executions. Additional details and the complete list of trigger words and their variants is provided in the Appendix.
Suppressing Triggers Generation via Certainty.
We reduce the reflection behaviors by suppressing the generation of triggers according to the certainty score. Specifically, we first calculate the suppression probability $p$ as follows,
$p=\max\left(0,\frac{C-\delta}{1-\delta}\right),$ (2)
where $\delta\in[0,1]$ is the confidence threshold. Then, with probability $p$, we set the logits of trigger tokens in $S_{\text{trigger}}$ to a large negative value, thereby effectively excluding them from sampling. Eq. (2) indicates that trigger suppression occurs only when $C>\delta$. Besides, higher $C$ values yield more frequent trigger suppression, as the LLM has exhibited high confidence in current responses, thereby requiring fewer subsequent reflection behaviors.
In summary, during inference, CGRS inserts checkpoints to probe intermediate answers for certainty calculation, then suppresses reflection tokens accordingly. This dynamic control mechanism reduces redundant reflection behaviors while preserving necessary corrections, enabling efficient reasoning. Moreover, CGRS is model-agnostic, requires no retraining or architectural modifications, and can be seamlessly integrated into existing autoregressive generation pipelines. The token prediction process of CGRS is shown in Algorithm 1.
4 Experiments
Method AIME24 AMC23 MATH500 GPQA-D AVG
[…]
Qwen3-4B
Vanilla 60.0 11449 - 91.7 7449 - 92.7 4796 - 53.5 8195 - 74.5 -
TALE 48.9 9727 15.0% 86.7 5107 31.4% 89.1 2657 44.6% 36.2 4938 39.7% 65.2 32.7%
NoThinking 24.4 4504 60.7% 70.0 1710 77.0% 84.9 988 79.4% 47.5 1471 82.1% 56.7 74.8%
Dynasor 54.3 9912 13.4% 86.7 6233 16.3% 90.1 3877 19.2% 50.5 4398 46.3% 70.4 23.8%
DEER 50.0 6873 40.0% 80.0 3231 56.6% 87.0 1854 61.3% 54.9 7033 14.2% 68.0 43.0%
CGRS (ours) 56.7 7893 31.1% 86.7 4351 41.6% 91.3 2704 43.6% 55.2 5229 36.2% 72.5 38.1%
Qwen3-8B
Vanilla 61.1 11924 - 89.4 7876 - 92.9 5104 - 57.7 9104 - 75.3 -
TALE 68.9 10942 8.2% 88.3 6872 12.7% 92.3 3885 23.9% 59.1 7113 21.9% 77.2 16.7%
NoThinking 30.0 5967 50.0% 72.5 2426 69.2% 87.1 1239 75.7% 54.2 1546 83.0% 61.0 69.5%
Dynasor 62.2 10174 14.7% 89.2 6457 18.0% 91.7 3841 24.7% 57.7 5965 34.5% 75.2 23.0%
DEER 45.6 7443 37.6% 79.2 3715 52.8% 88.7 1935 62.1% 59.3 7837 13.9% 68.2 41.6%
CGRS (ours) 61.1 8792 26.3% 89.2 5595 29.0% 93.3 3507 31.3% 59.8 6302 30.8% 75.9 29.3%
Qwen3-14B
Vanilla 68.9 11316 - 93.3 7190 - 94.1 4551 - 64.0 7411 - 80.1 -
TALE 71.1 10860 6.4% 92.5 5951 17.2% 93.7 3389 25.5% 63.8 6091 17.8% 80.3 16.2%
NoThinking 27.8 3689 67.4% 77.5 $16$ 77.5% 87.0 853 81.3% 56.9 1268 82.9% 62.3 77.3%
Dynasor 65.6 9775 13.6% 90.0 6030 16.1% 84.4 3667 19.4% 64.3 5775 22.1% 76.1 17.8%
DEER 56.7 6755 40.3% 90.8 4079 43.3% 91.7 1956 57.0% 58.2 6338 14.5% 74.4 38.8%
CGRS (ours) 70.0 8662 23.5% 93.3 5076 29.4% 94.5 3235 28.9% 65.2 5953 19.7% 80.8 25.4%
DeepSeek-R1-Distill-Qwen-7B
Vanilla 52.2 10662 - 87.5 5861 - 91.3 3787 - 36.2 7191 - 66.8 -
TALE 48.9 9727 8.8% 86.7 5107 19.9% 89.1 2657 29.8% 36.2 4938 31.3% 65.2 20.7%
NoThinking 32.2 6680 37.3% 75.8 2499 57.4% 80.9 1173 65.4% 37.9 1312 81.8% 56.7 61.4%
Dynasor 47.8 8334 21.8% 84.2 5201 11.3% 81.8 2070 45.3% 22.2 561 92.2% 59.0 42.7%
DEER 47.8 9288 12.9% 88.3 4670 20.3% 89.6 2272 40.0% 33.1 6457 10.2% 64.7 20.9%
CGRS (ours) 52.2 7597 28.7% 88.3 3406 41.9% 87.6 1867 50.7% 32.8 3876 46.1% 65.2 41.9%
DeepSeek-R1-Distill-Llama-8B
Vanilla 37.7 11898 - 84.2 6374 - 85.7 4087 - 39.2 8096 - 61.7 -
TALE 40.0 11141 6.4% 85.0 5915 7.2% 84.0 3541 13.4% 46.1 5573 31.2% 63.8 14.5%
NoThinking 40.0 11242 5.5% 82.5 5796 9.1% 83.3 2405 41.2% 36.2 6638 18.0% 60.5 18.4%
Dynasor 37.7 10368 12.9% 84.2 5681 10.9% 85.0 3585 12.3% 31.8 3095 61.8% 59.7 24.4%
DEER 42.2 9778 17.8% 80.8 5480 14.0% 82.3 2722 33.4% 41.8 7434 8.2% 56.4 35.7%
CGRS (ours) 47.8 9536 19.9% 86.7 4899 23.1% 84.7 3254 20.4% 39.6 7221 10.8% 64.7 18.5%
QwQ-32B
Vanilla 71.1 11026 - 89.1 7210 - 94.2 4216 - 66.4 7269 - 80.2 -
TALE 61.1 10888 1.3% 90.8 6522 9.5% 94.0 3533 16.2% 65.5 6482 10.8% 77.8 9.5%
NoThinking 62.2 11688 -6.0% 88.3 7493 -3.9% 94.2 4276 -1.4% 65.2 7604 -4.6% 77.5 -4.0%
Dynasor 64.4 9733 11.7% 90.0 7185 0.3% 94.0 4156 1.4% 41.9 2478 65.9% 72.6 19.9%
DEER 65.6 10015 9.2% 92.5 6324 12.3% 94.1 3359 20.3% 66.5 6453 11.2% 79.7 13.3%
CGRS (ours) 68.9 8202 25.6% 93.3 4771 33.8% 94.2 2810 33.3% 67.0 5141 29.3% 80.8 30.5%
Table 1: Comparison across models of different scales and multiple methods. Each experiment is repeated three times and the average results are reported. “Acc” (%) and “Len” (in tokens) denote the accuracy and response length, respectively. “LR” is the average length reduction ratio relative to Vanilla. $\uparrow$ ($\downarrow$) indicates that the higher (lower) the result, the better the performance. The best results (except for Vanilla) are highlighted in bold.
In this section, we evaluate the proposed CGRS method on multiple benchmark datasets and models to demonstrate its effectiveness in enabling efficient reasoning.
4.1 Experimental Setup
Benchmark Datasets.
We evaluate on three mathematical reasoning benchmarks: AIME24 (MAA Committees 2025) that contains $30$ high-difficulty problems from the 2024 American Invitational Mathematics Examination, AMC23 (AI-MO 2024) that includes $40$ problems from the 2023 American Mathematics Competitions, and MATH500 (Lightman et al. 2023) that has $500$ multi-step problems covering algebra, geometry, and probability, curated by OpenAI, and one scientific reasoning benchmark: GPQA Diamond (abbreviated as GPQA-D) (Rein et al. 2024), a dataset of $198$ multiple-choice questions on biology, chemistry, and physics, at the post-graduate level.
Models.
We conduct experiments on the a series of open-source models spanning different architectures and scales, including DeepSeek-R1-Distill series of models (Qwen-7B and Llama-8B) (Guo et al. 2025), QwQ-32B (Team 2025), and the Qwen3 family (4B, 8B, 14B, and 32B) (Yang et al. 2025a).
Implementation Details.
All experiments are conducted using the open-source vLLM framework (Kwon et al. 2023) to ensure high-throughput and memory-efficient inference. All decoding uses a temperature of $0.6$ and top-p of $0.95$. Each experiment is repeated three times and the average results are reported. The reflection suppression threshold $\delta$ in our CGRS method is set to $0.9$.
Baselines.
The proposed CGRS method is compared with three types of baselines: (i) Vanilla that performs standard decoding without any intervention, (ii) prompt-guided methods, including NoThinking (Ma et al. 2025a) that prompts the model to bypass intermediate reasoning entirely and generate the final answer directly, and TALE (Han et al. 2025) that prompts the model to solve the problem within token budgets (in our experiments, we set this budget based on the actual token length generated by CGRS), and (iii) decoding-manipulation methods, including Dynasor (Fu et al. 2024) that periodically requests intermediate answers at fixed token intervals and exits early if multiple consecutive answers match and DEER (Yang et al. 2025b) dynamically truncates chain-of-thought generation by detecting high-confidence intermediate answers at transition cues such as “Wait”.
Evaluation Metrics.
We evaluate performance using three metrics: (i) pass@1 accuracy (Acc) that is the proportion of problems correctly solved on the first attempt, (ii) average output token length (Len) that serves as a proxy for reasoning cost during inference, and (iii) the average length reduction ratio (LR) that measures the percentage decrease in output token length compared to the Vanilla method, with higher values indicating higher compression.
4.2 Comparison with State-of-the-art Methods
In Table 1, we compare the proposed CGRS method with previous efficient reasoning methods in terms of accuracy and output length. The evaluation covers diverse model scales and architectures, including DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, QwQ-32B, and Qwen3-4/8/14/32B across four reasoning benchmark datasets (AIME24, AMC23, MATH500, and GPQA-D). The results for Qwen3-32B are provided separately in the Appendix.
As can be seen, CGRS consistently achieves the optimal balance between average length reduction ratio and average accuracy preservation across diverse models, demonstrating its effectiveness for efficient reasoning. For example, on the Qwen3-14B model, CGRS reduces output length by $25.4\%$ while maintaining accuracy. In contrast, while baseline methods achieve greater length reduction ($77.3\%$ for NoThinking and $38.8\%$ for DEER), they suffer significant accuracy degradation ($17.8\%$ and $5.7\%$ drops, respectively). TALE maintains comparable accuracy but achieves less length reduction than CGRS ($16.2\%$ vs. $25.4\%$).
Moreover, the proposed CGRS method demonstrates reliable effectiveness in every test case, consistently achieving $18.5\%$ to $41.9\%$ length reduction with negligible accuracy drop (up to $2\%$), while the baseline methods often perform unsatisfactorily. For example, NoThinking prompts models to bypass slow thinking processes, which yields high compression rates but severely degrades accuracy (e.g., $65.4\%$ length reduction and $9.4\%$ accuracy drop on MATH500 with DeepSeek-R1-Distill-Qwen-7B). Early-exit approaches like Dynasor and DEER also sometimes underperform, e.g., on AMC23 with DeepSeek-R1-Distill-Llama-8B, they only achieve $10.9\%$ and $14.0\%$ compression, respectively, with DEER additionally reducing accuracy. Remarkably, in the same test case, CGRS attains $23.1\%$ length reduction while slightly improving accuracy.
Notably, on the QwQ-32B model, the baseline methods (TALE, NoThinking, Dynasor, and DEER) achieve only minor reduction ratios across all four benchmark datasets. This limitation occurs because these methods rely on hard-coded reflection boundaries that assume reasoning terminates at the </think> token, while QwQ-32B often continues reflecting beyond this token (Yang et al. 2025b). In contrast, the proposed CGRS method dynamically regulates reflection triggers based on confidence scores, which is independent on the end-of-think token, enabling CGRS to maintain strong compression rates ($30.5\%$) while preserving accuracy in this case.
In summary, these results demonstrate the superior performance of the proposed CGRS method over all baseline approaches across various models and datasets, indicating its effectiveness for efficient reasoning in LRLMs.
4.3 Effectiveness of Reflection Suppression
As introduced in Section 3, CGRS proactively suppresses redundant reflection behaviors to alleviate the overthinking problem. To evaluate the effectiveness of this mechanism, in this section, we analyze two key metrics: (i) the frequency of reflection triggers (specifically the words “Wait”, “But”, “Alternatively”, and “Hmm”), and (ii) the answer length distribution, comparing CGRS against the Vanilla baseline. The experiments employ the DeepSeek-R1-Distill-Qwen-7B model on the AIME24 benchmark, following the same experimental setup described in Section 4.1.
Figure 2: Comparison of Vanilla and CGRS methods in terms of the frequency of reflection triggers (left) and the answer length distribution (right).
The results are presented in Figure 2. Compared to the Vanilla baseline, CGRS demonstrates: (i) a significant reduction in reflection trigger words (particularly “Wait” and “But”), and (ii) a more concentrated and significantly shorter length distribution. These findings clearly indicate effective suppression of reflection behaviors. When combined with the findings in Table 1 showing that CGRS achieves 28.7% length reduction without accuracy decrease on AIME24 using the DeepSeek-R1-Distill-Qwen-7B model, these results demonstrate that CGRS effectively alleviate the overthinking problem and enable efficient reasoning.
4.4 Effectiveness of Certainty-Guided Probability
As described in Section 3, CGRS dynamically suppresses the LRLM’s tendency to generate reflection triggers based on the certainty of its current response. In this section, we evaluate the effectiveness of this certainty-guided suppression probability (Eq. (2)) by comparing it against fixed suppression probabilities: $p=0$ (the Vanilla baseline described in Section 4.1), $0.25$, $0.5$, and $1$ (where no reflection triggers are sampled during inference). We conduct experiments on the AMC23 dataset using the DeepSeek-R1-Distill-Qwen-7B model, following the same setup as in Section 4.1.
[…]
$p=0$ (the Vanilla baseline) 87.5 5861
$p=0.25$ 81.7 3729
$p=0.5$ 80.0 3266
$p=1.0$ 76.7 2373
Certainty-Guided (Eq. (2)) 88.3 3406
Table 2: Comparison of certainty-guided suppression probability (Eq. (2)) with fixed probabilities.
Figure 3: Comparison of the responses from Vanilla and the proposed CGRS (with $\delta$ = 0.9) methods on an example from the MATH500 dataset using the DeepSeek-R1-Distill-Qwen-7B model.
The results are shown in Table 2. As can be seen, as $p$ increases, the response length decreases significantly while the accuracy also drops considerably. This indicates that important reflection behaviors are being suppressed in the slow thinking process. In contrast, our certainty-guided method achieves a 41.9% reduction in length without compromising accuracy, demonstrating that this certainty-based control effectively suppresses unproductive reflections while preserving necessary ones.
4.5 Case Study
In Figure 3, we present a representative case from the MATH 500 dataset, solved by the DeepSeek-R1-Distill-7B model, to demonstrate CGRS’s mechanism and advantages.
During the initial reasoning phase, the model executes routine computations. Since the certainty scores for these intermediate steps remain below a predefined threshold, CGRS remains inactive, permitting the model to freely explore and revise its reasoning path—similar to standard autoregressive inference.
The turning point occurs when a high certainty score of $0.926$ is detected, triggering the CGRS mechanism. In the Vanilla baseline, the model performs redundant reflection even after reaching a correct intermediate answer—re-examining previous steps (e.g., common ratio, exponentiation, and multiplication in this case). This unnecessary verification increases computational overhead without enhancing accuracy. In contrast, CGRS promotes efficient and goal-directed behavior. It suppresses reflection-inducing tokens, preventing re-evaluation of outputs it considers reliable. Instead, the model proceeds directly to the next logical operation (e.g., fraction simplification). This targeted intervention ensures high-confidence reasoning proceeds without unnecessary detours, significantly reducing both token count and inference time.
This case study offers qualitative evidence that CGRS effectively alleviate overthinking by utilizing internal certainty estimates to suppress redundant reasoning steps.
5 Conslusion
In this paper, we propose Certainty-Guided Reflection Suppression (CGRS), a novel approach to address the overthinking problem in Large Reasoning Language Models (LRLMs) and enable efficient reasoning. CGRS is a lightweight, certainty-guided decoding strategy that dynamically suppresses reflection triggers when the model exhibits high confidence in its current reasoning trajectory. Our method is model-agnostic, requires no retraining or architectural modifications, and can be seamlessly integrated into existing autoregressive generation pipelines. Extensive experiments on open-source LRLMs across four reasoning benchmarks demonstrate that CGRS achieves token usage reductions of up to $41.9\%$ while preserving answer accuracy. These results establish that reflective behaviors in LRLMs can be effectively modulated during inference through certainty-aware decoding. Our work advances the development of efficient reasoning systems by demonstrating how model behavior can be aligned with internal certainty signals, providing a practical pathway toward more scalable LLM deployments.

### v2

@@
 Our approach is model-agnostic, requires no retraining or architectural modifications, and can be integrated seamlessly with existing autoregressive generation pipelines.
-Extensive experiments across four reasoning benchmarks (i.e., AIME24, AMC23, MATH500, and GPQA-D) demonstrate CGRS’s effectiveness: it reduces token usage by an average of $18.5\%$ to $41.9\%$ while preserving accuracy.
-It also achieves the optimal balance between length reduction and performance compared to state-of-the-art baselines.
+Extensive experiments across four reasoning benchmarks (i.e., AIME24, AMC23, MATH500, and GPQA-D) demonstrate CGRS’s effectiveness: it reduces token usage by an average of $18.5\%$ to $41.9\%$ while preserving accuracy and also achieves the optimal balance between length reduction and performance compared to state-of-the-art baselines.
 These results hold consistently across model architectures (e.g., DeepSeek-R1-Distill series, QwQ-32B, and Qwen3 family) and scales (4B to 32B parameters), highlighting CGRS’s practical value for efficient reasoning.
@@
 1 Introduction
-Large Reasoning Language Models (LRLMs), including OpenAI’s o1/o3 (OpenAI 2024, 2025) and DeepSeek-R1 (Guo et al.
-2025), have demonstrated remarkable performance on demanding benchmarks, especially in advanced mathematics and program synthesis.
-A key factor behind their success is their slow-thinking capability, which combines step-by-step deduction with complex reflection behaviors, including backtracking, exploring alternative strategies, and self-verification of results  (Zhu and Li 2025).
-Although these reflection behaviors enable models to self-correct and obtain accurate answers, they often suffer from a significant overthinking problem (Cuadron et al.
-2025; Sui et al.
-2025; Fan et al.
-2025), where LRLMs persist in reasoning even after arriving at correct solutions.
-This phenomenon leads to substantial increases in token consumption, higher inference costs, and a degraded user experience due to unnecessary delays (Sun et al.
-2025; Wang et al.
-2025b).
-In extreme cases, excessively long responses may exceed context window limits, truncating critical information and compromising accuracy (Li et al.
-2025a).
+Large Reasoning Language Models (LRLMs), including OpenAI’s o1/o3 (o1; o3) and DeepSeek-R1 (guo2025deepseekr1), have demonstrated remarkable performance on demanding benchmarks, especially in advanced mathematics and program synthesis.
+A key factor behind their success is their slow-thinking capability, which combines step-by-step deduction with complex reflection behaviors, including backtracking, exploring alternative strategies, and self-verification of results  (zhu2025conciseadaptivethinkinglarge).
+Although these reflection behaviors enable models to self-correct and obtain accurate answers, they often suffer from a significant overthinking problem (cuadron2025dangeroverthinkingexaminingreasoningaction; sui2025stopoverthinkingsurveyefficient; fan2025missingpremiseexacerbatesoverthinking), where LRLMs persistently continue reasoning even after arriving at correct solutions.
+This phenomenon leads to substantial increases in token consumption, higher inference costs, and degraded user experience due to unnecessary delays (sun2025invisibletokensvisiblebills; wang2025harnessingreasoningeconomysurvey).
+In extreme cases, excessively prolonged responses may exceed context window limits, resulting in truncated critical information and compromised accuracy (li2025thinklesstrainingfreeinferenceefficientmethod).
 Addressing this overthinking problem and developing efficient reasoning mechanisms, therefore represents a crucial challenge for improving LRLM performance and practicality.
-To address this challenge, training-free methods have recently emerged as lightweight solutions that mitigate overthinking without requiring additional training (Zhu and Li 2025).
-For instance, prompt-guided approaches like TALE (Han et al.
-2025) prompts model to respond within predefined token budgets through carefully designed instructions.
+To address this challenge, training-free methods have recently emerged as lightweight solutions that mitigate overthinking without requiring additional training (zhu2025conciseadaptivethinkinglarge).
+For instance, prompt-guided approaches like TALE (han-etal-2025-token) prompts model responses within predefined token budgets through carefully designed instructions.
 While straightforward, these methods rely heavily on the model’s innate instruction-following capabilities.
-Decoding-manipulation techniques, such as Dynasor (Fu et al.
-2024) and DEER (Yang et al.
-2025b) dynamically adjust the decoding process to eliminate redundancy by terminating generation when the model reaches sufficient confidence.
+Decoding-manipulation techniques, such as Dynasor (fu2025efficientlyscalingllmreasoning) and DEER (yang2025dynamicearlyexitreasoning), dynamically adjust the decoding process to eliminate redundancy by terminating generation when the model reaches sufficient confidence.
 However, their effectiveness remains highly sensitive to the design of early-exit conditions.
@@
 We propose Certainty-Guided Reflection Suppression (CGRS), a novel method that mitigates overthinking in LRLMs while preserving reasoning accuracy.
-Previous work has demonstrated that reflection behaviors are typically signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (Muennighoff et al.
-2025; Zhang et al.
-2025; Ding et al.
-2025).
-Our CGRS approach proactively suppresses the model’s generation of these reflection triggers when it exhibits high certainty in its current response, thereby preventing redundant reflection cycles, effectively alleviating overthinking, and preserving reasoning quality.
+Previous work has demonstrated that reflection behaviors are typically signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (muennighoff2025s1; zhang2025alphaonereasoningmodelsthinking; ding2025thinkingtokenshelptrap).
+Our CGRS approach proactively suppresses the model’s generation of these reflection triggers when it exhibits high certainty in its current response, thereby preventing redundant reflection cycles and effectively alleviating overthinking while preserving reasoning quality.
 […]
-We evaluate CGRS across diverse open-source LRLMs spanning multiple architectures and scales: the DeepSeek-R1-Distill series (Qwen-7B and Llama-8B) (Guo et al.
-2025), QwQ-32B (Team 2025), and the Qwen3 family (4B, 8B, 14B, and 32B) (Yang et al.
-2025a).
-Our evaluation covers four reasoning benchmarks, including mathematical tasks (AIME24 (MAA Committees 2025), AMC23 (AI-MO 2024), and MATH500 (Lightman et al.
-2023)) and scientific reasoning (GPQA Diamond (Rein et al.
-2024)).
+We evaluate CGRS across diverse open-source LRLMs spanning multiple architectures and scales: the DeepSeek-R1-Distill series (Qwen-7B and Llama-8B) (guo2025deepseekr1), QwQ-32B (qwq32b), and the Qwen3 family (4B, 8B, 14B, and 32B) (qwen3).
+Our evaluation covers four reasoning benchmarks, including mathematical tasks (AIME24 (aime), AMC23 (amc23), and MATH500 (math500)) and scientific reasoning (GPQA Diamond (gpqa)).
 As shown in Fig.
@@
 2 Related Work
-LRLMs have significantly advanced long Chain-of-Thought (CoT) reasoning through supervised fine-tuning and reinforcement learning (Pan et al.
-2025).
-However, prolonged CoT often introduces redundant steps that increase computational costs, a phenomenon known as overthinking (Cuadron et al.
-2025; Sui et al.
-2025; Fan et al.
-2025).
-To address this issue, many works focus on compressing verbose CoTs into more concise reasoning traces (Sui et al.
-2025; Qu et al.
-2025; Feng et al.
-2025; Wang et al.
-2025b; Zhu and Li 2025), which can be broadly categorized into two approaches: training-based methods (Arora and Zanette 2025; Aggarwal and Welleck 2025; Wang et al.
-2025a; Ma et al.
-2025b; Munkhbat et al.
-2025; Chen et al.
-2025b; Zhu et al.
-2025) and training-free methods (Ma et al.
-2025a; Yu, Yu, and Wang 2025; Han et al.
-2025; Yang et al.
-2025c; Team et al.
-2025; Wu et al.
-2025; Fu et al.
-2024; Jiang et al.
-2025; Yang et al.
-2025b; Li et al.
-2025b; Aytes, Baek, and Hwang 2025; Yang et al.
-2025c).
+LRLMs have significantly advanced long Chain-of-Thought (CoT) reasoning through supervised fine-tuning and reinforcement learning (Pan2025ASO).
+However, prolonged CoT often introduces redundant steps that increase computational costs, a phenomenon known as overthinking (cuadron2025dangeroverthinkingexaminingreasoningaction; sui2025stopoverthinkingsurveyefficient; fan2025missingpremiseexacerbatesoverthinking).
+To address this issue, many works focus on compressing verbose CoTs into more concise reasoning traces (sui2025stopoverthinkingsurveyefficient; qu2025surveyefficientreasoninglarge; feng2025efficientreasoningmodelssurvey; wang2025harnessingreasoningeconomysurvey; zhu2025conciseadaptivethinkinglarge), which can be broadly categorized into two approaches: training-based methods (arora2025traininglanguagemodelsreason; aggarwal2025l1controllinglongreasoning; wang2025learningthinkinformationtheoreticreinforcement; ma2025cotvalvelengthcompressiblechainofthoughttuning; munkhbat2025selftrainingelicitsconcisereasoning; chen2025reasoninglanguagecomprehensivesurvey; zhu2025surveylatentreasoning) and training-free methods (ma2025reasoningmodelseffectivethinking; yu2025premisescalablestrategicprompt; han-etal-2025-token; yang2025speculativethinkingenhancingsmallmodel; kimiteam2025kimik15scalingreinforcement; wu2025unlockingefficientlongtoshortllm; fu2025efficientlyscalingllmreasoning; jiang2025flashthinkearlyexitmethod; yang2025dynamicearlyexitreasoning; li2025steeringllmthinkingbudget; aytes2025sketchofthoughtefficientllmreasoning; yang2025speculativethinkingenhancingsmallmodel).
 While effective, training-based methods often require retraining, careful data curation, and may compromise the general capabilities of the pretrained model.
@@
 Prompt-guided Methods.
-This type of method (Ma et al.
-2025a; Yu, Yu, and Wang 2025; Han et al.
-2025) compress lengthy CoT reasoning through prompt engineering.
-For instance, TALE (Han et al.
-2025) simply prompts the model to solve the problem within token budgets.
+This type of method (ma2025reasoningmodelseffectivethinking; yu2025premisescalablestrategicprompt; han-etal-2025-token) compress lengthy CoT reasoning through prompt engineering.
+For instance, TALE (han-etal-2025-token) simply prompts the model to solve the problem within token budgets.
 While simple, these methods heavily depend on the model’s inherent instruction-following capability.
 Pipeline-based Methods.
-This category of methods (Ong et al.
-2025; Aytes, Baek, and Hwang 2025; Yang et al.
-2025c) distribute queries or reasoning stages across multiple LRLMs, leveraging the efficiency of small models and the advanced reasoning of larger ones.
+This category of methods (ong2025routellmlearningroutellms; aytes2025sketchofthoughtefficientllmreasoning; yang2025speculativethinkingenhancingsmallmodel) distribute queries or reasoning stages across multiple LRLMs, leveraging the efficiency of small models and the advanced reasoning of larger ones.
 However, these approaches introduce deployment complexity and additional inference overhead from auxiliary modules.
 Model Merging.
-Such methods (Team et al.
-2025; Wu et al.
-2025) combine slow- and fast-thinking models via weight averaging to enable long-to-short reasoning.
-While effective for medium-sized models, this approach underperforms at extreme scales (very large or small models) and lacks precise control over reasoning depth (Zhu and Li 2025).
+Such methods (kimiteam2025kimik15scalingreinforcement; wu2025unlockingefficientlongtoshortllm) combine slow- and fast-thinking models via weight averaging to enable long-to-short reasoning.
+While effective for medium-sized models, this approach underperforms at extreme scales (very large or small models) and lacks precise control over reasoning depth (zhu2025conciseadaptivethinkinglarge).
 Decoding-manipulation Methods.
-This type of method (Fu et al.
-2024; Jiang et al.
-2025; Yang et al.
-2025b; Li et al.
-2025b) dynamically adjusts the decoding process to eliminate redundancy.
-For example, Dynasor (Fu et al.
-2024) and DEER (Yang et al.
-2025b) terminate decoding early when the model reaches sufficient confidence, though their performance depends critically on the exit condition design.
+This type of method (fu2025efficientlyscalingllmreasoning; jiang2025flashthinkearlyexitmethod; yang2025dynamicearlyexitreasoning; li2025steeringllmthinkingbudget) dynamically adjusts the decoding process to eliminate redundancy.
+For example, Dynasor (fu2025efficientlyscalingllmreasoning) and DEER (yang2025dynamicearlyexitreasoning) terminate decoding early when the model reaches sufficient confidence, though their performance depends critically on the exit condition design.
 In this work, we focus on decoding manipulation for efficient reasoning.
@@
 LRLMs organize their output using thinking delimiters (i.e., <think> and </think>), splitting the response into two key parts: slow thinking and conclusion.
-During slow thinking, LRLMs perform detailed, step-by-step reasoning before summarizing their thought process and delivering the final answer in the conclusion (Zhang et al.
-2025).
+During slow thinking, LRLMs perform detailed, step-by-step reasoning before summarizing their thought process and delivering the final answer in the conclusion (zhang2025alphaonereasoningmodelsthinking).
 Within slow thinking, models engage in complex reflection behaviors, including backtracking, exploring alternative approaches, and verifying results.
-These behaviors are often signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (Muennighoff et al.
-2025; Zhang et al.
-2025; Ding et al.
-2025).
+These behaviors are often signaled by specific reflection triggers, i.e., keywords like “Wait”, “Alternatively”, “But”, and “Hmm” (muennighoff2025s1; zhang2025alphaonereasoningmodelsthinking; ding2025thinkingtokenshelptrap).
 Notably, these reflection behaviors cause models to continue reasoning even after reaching an initial answer.
 They may perform extra validation steps or switch reasoning strategies in later stages, either to double-check previous results or explore different paths.
-However, this can sometimes lead to unproductive overthinking, where the model enters unnecessary reasoning loops, repeatedly validating or re-evaluating without meaningful progress (Chen et al.
-2025a).
+However, this can sometimes lead to unproductive overthinking, where the model enters unnecessary reasoning loops, repeatedly validating or re-evaluating without meaningful progress (chen2025think23overthinkingo1like).
 To address this, our CGRS method proactively suppresses the model’s tendency to generate reflection triggers when it exhibits high certainty in its current response.
@@
 3.2 Certainty Estimation
-To estimate the necessity of reflective behaviors during inference, we introduce a certainty score that quantifies the LLM’s confidence in its currently generated response.
 […]
 $C=1-\left(\frac{\frac{1}{n}\sum_{i=1}^{n}\mathcal{H}({\mathbf{p}}_{{\mathbf{a}}_{i}})}{\log(|{\mathbf{V}}|)}\right),$ (1)
-where $n=|{\mathbf{a}}|$ is the token length of the probed result, ${\mathbf{p}}_{{\mathbf{a}}_{i}}={\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t},{\mathbf{c}},{\mathbf{a}}_{<i})$ represents the probability distribution of ${\mathbf{a}}_{i}$, $\mathcal{H}({\mathbf{p}})=-{\mathbf{p}}^{\top}\log({\mathbf{p}})$ denotes the token entropy, and $|{\mathbf{V}}|$ is the total number of tokens in the LLM’s vocabulary.
+where $n=|{\mathbf{a}}|$ is the token length of the probed results, ${\mathbf{p}}_{{\mathbf{a}}_{i}}={\bm{\pi}}(\cdot|{\mathbf{q}},{\mathbf{o}}_{<t},{\mathbf{c}},{\mathbf{a}}_{<i})$ represents the probability distribution of ${\mathbf{a}}_{i}$, $\mathcal{H}({\mathbf{p}})=-{\mathbf{p}}^{\top}\log({\mathbf{p}})$ denotes the token entropy, and $|{\mathbf{V}}|$ is the total number of tokens in the LLM’s vocabulary.
 The term $\log(|{\mathbf{V}}|)$ represents the maximum possible entropy (i.e., when the probability distribution is uniform across all tokens), serving as a normalization factor.
@@
 Reflection Triggers.
-Following prior work (Yang et al.
-2025c; Fu et al.
-2024; Zhang et al.
-2025), we consider three trigger categories: (i) core hesitation and transition words (e.g., “But” and “Wait”), (ii) alternative proposal markers (e.g., “Alternatively”), and (iii) colloquial contemplation cues (e.g., “Hmm”), all of which signal potential shifts in the LLM’s reasoning mode.
+Following prior work (yang2025speculativethinkingenhancingsmallmodel; fu2025efficientlyscalingllmreasoning; zhang2025alphaonereasoningmodelsthinking), we consider three trigger categories: (i) core hesitation and transition words (e.g., “But” and “Wait”), (ii) alternative proposal markers (e.g., “Alternatively”), and (iii) colloquial contemplation cues (e.g., “Hmm”), all of which signal potential shifts in the LLM’s reasoning mode.
 Since a single trigger word has multiple natural language variants (e.g., “wait” as a lowercase variant of “Wait”) that map to different tokens in the vocabulary, we construct a comprehensive trigger token set $S_{\text{trigger}}$ through frequency analysis.
 Specifically, we identify all possible variants of each reflection trigger in the tokenizer’s vocabulary and then analyze their generation frequencies in reasoning traces.
-This analysis was performed using reasoning traces from the DeepSeek-R1-Distill-Qwen-7B model (Guo et al.
-2025) on both the AIME24 (MAA Committees 2025) and AMC23 (AI-MO 2024) benchmarks.
+This analysis was performed using reasoning traces from the DeepSeek-R1-Distill-Qwen-7B model (guo2025deepseekr1) on both the AIME24 (aime) and AMC23 (amc23) benchmarks.
 To ensure robustness, we conducted four independent inference runs per dataset, aggregating statistics across all executions.
-Additional details and the complete list of trigger words and their variants is provided in the Appendix.
+Additional details and the complete list of trigger words and their variants are available in Appendix.
 Suppressing Triggers Generation via Certainty.
@@
 CGRS (ours) 70.0 8662 23.5% 93.3 5076 29.4% 94.5 3235 28.9% 65.2 5953 19.7% 80.8 25.4%
+Qwen3-32B
+Vanilla 67.8 11022 - 95.8 6794 - 93.9 4473 - 71.4 7062 - 82.2 -
+TALE 67.8 10688 3.0% 93.3 6533 3.8% 93.6 3857 13.8% 67.0 5916 16.2% 80.4 9.2%
+NoThinking 41.1 5635 48.9% 75.0 2221 67.3% 87.0 1054 76.4% 56.6 917 87.0% 64.9 69.9%
+Dynasor 64.4 9518 13.6% 92.5 5521 18.7% 85.2 3486 22.1% 58.1 3161 55.2% 75.0 27.4%
+DEER 68.9 8697 21.1% 88.2 4878 28.2% 93.2 2533 43.4% 69.6 6310 10.6% 80.0 25.8%
+CGRS (ours) 65.6 8128 26.3% 94.2 4766 29.8% 93.1 2993 33.1% 64.0 4535 35.8% 79.2 31.2%
 DeepSeek-R1-Distill-Qwen-7B
@@
 Benchmark Datasets.
-We evaluate on three mathematical reasoning benchmarks: AIME24 (MAA Committees 2025) that contains $30$ high-difficulty problems from the 2024 American Invitational Mathematics Examination, AMC23 (AI-MO 2024) that includes $40$ problems from the 2023 American Mathematics Competitions, and MATH500 (Lightman et al.
-2023) that has $500$ multi-step problems covering algebra, geometry, and probability, curated by OpenAI, and one scientific reasoning benchmark: GPQA Diamond (abbreviated as GPQA-D) (Rein et al.
-2024), a dataset of $198$ multiple-choice questions on biology, chemistry, and physics, at the post-graduate level.
+We evaluate on three mathematical reasoning benchmarks: AIME24 (aime) that contains $30$ high-difficulty problems from the 2024 American Invitational Mathematics Examination, AMC23 (amc23) that includes $40$ problems from the 2023 American Mathematics Competitions, and MATH500 (math500) that has $500$ multi-step problems covering algebra, geometry, and probability, curated by OpenAI, and one scientific reasoning benchmark: GPQA Diamond (abbreviated as GPQA-D) (gpqa), a dataset of $198$ multiple-choice questions on biology, chemistry, and physics, at the post-graduate level.
 Models.
-We conduct experiments on the a series of open-source models spanning different architectures and scales, including DeepSeek-R1-Distill series of models (Qwen-7B and Llama-8B) (Guo et al.
-2025), QwQ-32B (Team 2025), and the Qwen3 family (4B, 8B, 14B, and 32B) (Yang et al.
-2025a).
+We conduct experiments on the a series of open-source models spanning different architectures and scales, including DeepSeek-R1-Distill series of models (Qwen-7B and Llama-8B) (guo2025deepseekr1), QwQ-32B (qwq32b), and the Qwen3 family (4B, 8B, 14B, and 32B) (qwen3).
 Implementation Details.
-All experiments are conducted using the open-source vLLM framework (Kwon et al.
-2023) to ensure high-throughput and memory-efficient inference.
-All decoding uses a temperature of $0.6$ and top-p of $0.95$.
+All experiments are conducted using the open-source vLLM framework (vllm) to ensure high-throughput and memory-efficient inference.
+All decoding use temperature $0.6$ and top-p $0.95$.
 Each experiment is repeated three times and the average results are reported.
@@
 Baselines.
-The proposed CGRS method is compared with three types of baselines: (i) Vanilla that performs standard decoding without any intervention, (ii) prompt-guided methods, including NoThinking (Ma et al.
-2025a) that prompts the model to bypass intermediate reasoning entirely and generate the final answer directly, and TALE (Han et al.
-2025) that prompts the model to solve the problem within token budgets (in our experiments, we set this budget based on the actual token length generated by CGRS), and (iii) decoding-manipulation methods, including Dynasor (Fu et al.
-2024) that periodically requests intermediate answers at fixed token intervals and exits early if multiple consecutive answers match and DEER (Yang et al.
-2025b) dynamically truncates chain-of-thought generation by detecting high-confidence intermediate answers at transition cues such as “Wait”.
+The proposed CGRS method is compared with three types of baselines: (i) Vanilla that performs standard decoding without any intervention, (ii) prompt-guided methods, including NoThinking (ma2025reasoningmodelseffectivethinking) that prompts the model to bypass intermediate reasoning entirely and generate the final answer directly, and TALE (han-etal-2025-token) that prompts the model to solve the problem within token budgets (in our experiments, we set this budget based on the actual token length generated by CGRS), and (iii) decoding-manipulation methods, including Dynasor (fu2025efficientlyscalingllmreasoning) that periodically requests intermediate answers at fixed token intervals and exits early if multiple consecutive answers match and DEER (yang2025dynamicearlyexitreasoning) dynamically truncates chain-of-thought generation by detecting high-confidence intermediate answers at transition cues such as “Wait”.
 Evaluation Metrics.
@@
 In Table 1, we compare the proposed CGRS method with previous efficient reasoning methods in terms of accuracy and output length.
-The evaluation covers diverse model scales and architectures, including DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, QwQ-32B, and Qwen3-4/8/14/32B across four reasoning benchmark datasets (AIME24, AMC23, MATH500, and GPQA-D).
-The results for Qwen3-32B are provided separately in the Appendix.
+The evaluation covers diverse model scales and architectures, including DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Llama-8B, QwQ-32B, and Qwen3-4/8/14/32B across four reasoning benchmark datasets (i.e., AIME24, AMC23, MATH500, and GPQA-D).
 As can be seen, CGRS consistently achieves the optimal balance between average length reduction ratio and average accuracy preservation across diverse models, demonstrating its effectiveness for efficient reasoning.
 For example, on the Qwen3-14B model, CGRS reduces output length by $25.4\%$ while maintaining accuracy.
-In contrast, while baseline methods achieve greater length reduction ($77.3\%$ for NoThinking and $38.8\%$ for DEER), they suffer significant accuracy degradation ($17.8\%$ and $5.7\%$ drops, respectively).
+In contrast, while baseline methods achieve greater length reduction ($77.3\%$ of NoThinking and $38.8\%$ of DEER), they suffer significant accuracy degradation ($17.8\%$ and $5.7\%$ drops respectively).
 TALE maintains comparable accuracy but achieves less length reduction than CGRS ($16.2\%$ vs.
 $25.4\%$).
-Moreover, the proposed CGRS method demonstrates reliable effectiveness in every test case, consistently achieving $18.5\%$ to $41.9\%$ length reduction with negligible accuracy drop (up to $2\%$), while the baseline methods often perform unsatisfactorily.
-For example, NoThinking prompts models to bypass slow thinking processes, which yields high compression rates but severely degrades accuracy (e.g., $65.4\%$ length reduction and $9.4\%$ accuracy drop on MATH500 with DeepSeek-R1-Distill-Qwen-7B).
+Moreover, the proposed CGRS method demonstrates reliable effectiveness in every test case, consistently achieving $18.5\%$ to $41.9\%$ length reduction with negligible accuracy drop (up to $3\%$), while the baseline methods often perform unsatisfactorily.
+For example, NoThinking prompts models to bypass slow thinking processes, which always yields high compression rates but severely degrades accuracy (e.g., $65.4\%$ length reduction and $9.4\%$ accuracy drop on MATH500 with DeepSeek-R1-Distill-Qwen-7B).
 Early-exit approaches like Dynasor and DEER also sometimes underperform, e.g., on AMC23 with DeepSeek-R1-Distill-Llama-8B, they only achieve $10.9\%$ and $14.0\%$ compression, respectively, with DEER additionally reducing accuracy.
@@
 Notably, on the QwQ-32B model, the baseline methods (TALE, NoThinking, Dynasor, and DEER) achieve only minor reduction ratios across all four benchmark datasets.
-This limitation occurs because these methods rely on hard-coded reflection boundaries that assume reasoning terminates at the </think> token, while QwQ-32B often continues reflecting beyond this token (Yang et al.
-2025b).
+This limitation occurs because these methods rely on hard-coded reflection boundaries that assume reasoning terminates at the </think> token, while QwQ-32B often continues reflection behaviors beyond this token (yang2025dynamicearlyexitreasoning).
 In contrast, the proposed CGRS method dynamically regulates reflection triggers based on confidence scores, which is independent on the end-of-think token, enabling CGRS to maintain strong compression rates ($30.5\%$) while preserving accuracy in this case.
@@
 Our work advances the development of efficient reasoning systems by demonstrating how model behavior can be aligned with internal certainty signals, providing a practical pathway toward more scalable LLM deployments.
+Acknowledgments
+This work is supported in part by the State Key Laboratory of General Artificial Intelligence.
+DH is supported by National Science Foundation of China (NSFC62376007) and Beijing Natural Science Foundation (Z250001).
+Appendix A Additional Details of Reflection Triggers Selection in Section 3.3
+As described in Section 3.3, we focus on four primary trigger words (“Wait”, “But”, “Alternatively”, and “Hmm”) in our experiments.
+For each trigger word, we identify all natural language variants present in the tokenizer’s vocabulary.
+For example, the keyword “Wait” has variants such as “.Wait,” “␣WaitForSeconds,” “.wait,” “.WaitFor,” “ĉwait,”.
+Then we analyze their generation frequencies in reasoning, in order to filter out tokens that are unlikely to appear (e.g., “.Wait”) during actual reasoning.
+This analysis uses reasoning traces from the DeepSeek-R1-Distill-Qwen7B model on both the AIME24 and AMC23 benchmarks.
+We examine two tokenizer types: LlamaTokenizerFast (for DeepSeek-R1-Distill-Llama-8B) and Qwen2Tokenizer (for DeepSeek-R1-Distill-Qwen-7B, the Qwen-3 family, and QwQ-32B).
+The complete list of filtered trigger words (including variants) and their corresponding token IDs are provided in Table 3.
+Appendix B Examples of Redundant Reflection Behaviors
+In this section, we present qualitative examples demonstrating redundant reflection behaviors in LRLM reasoning.
+We analyze responses from the DeepSeek-R1-Distill-Qwen-7B model on a complex combinatorial problem in the AIME24 dataset.
+As shown in Table 5, the model initially identifies a correct or promising approach but subsequently enters a prolonged reflective phase.
+This phase is characterized by ineffective self-correction, repeated attempts at previously explored approaches, and overthinking trivial details—often initiated by reflection triggers (like “Wait”, “But” and “Alternatively”).
+Despite the extended reasoning, these additional steps fail to improve answer quality and, in some cases, introduce new errors or overwrite correct logic.
+These failure cases underscore the need for dynamic reflection control during inference.
+By identifying the onset of these inefficient reflective loops, CGRS helps maintain concise and effective reasoning without suppressing productive analysis.
+Trigger Words Qwen Token ID Llama Token ID
+Wait 14190 14524
+␣Wait 13824 14144
+wait 11489 11748
+␣wait 3783 3868
+But 3983 4071
+␣But 1988 2030
+but 8088 8248
+␣but 714 719
+Alternatively 38478 93114
+␣Alternatively 38478 39578
+Alternative 75763 76863
+␣Alternative 41109 42209
+Hmm 80022 81122
+␣Hmm 88190 89290
+Table 3: Reflection trigger words and their corresponding token IDs in Qwen2Tokenizer and LlamaTokenizerFast tokenizers. ␣ denotes a space character.
+Appendix C Ablation on Threshold $\delta$
+The results are shown in Table 4.
+As $\delta$ decreases, the model becomes less strict in suppressing uncertain reasoning steps, leading to longer responses but lower accuracy.
+Specifically, when $\delta$ drops from 0.9 to 0.1, accuracy decreases from 88.3% to 72.5%, while the length reduction (LR) weakens from 41.9% to 53.7%.
+This demonstrates the trade-off controlled by $\delta$: higher thresholds enforce more confident reasoning and concise outputs, whereas lower thresholds allow more exploratory but error-prone reasoning steps.
+[…]
+0.9 88.3% 41.9
+0.5 76.2% 48.5
+0.1 72.5% 53.7
+Table 4: Effect of threshold $\delta$ on accuracy and length reduction (LR) on AMC23 and R1-Distill-Qwen-7B.
+Reflection Examples Analysis
+Case 1: Failed self-correction.
+“Therefore, $r(S)$ is $S^{\prime}$ which is equivalent to $S\cap r^{-1}(S^{\prime})=S$.
+[Correct] Wait, maybe that’s not the right way.
+Alternatively, since $r$ is a permutation, $r(S)$ is just another set of the same size as $S$.
+So, if $S^{\prime}$ is a subset of $S$, then $r(S)$ must be a subset of $S^{\prime}$ which would require that $S^{\prime}$ is a subset of $r^{-1}(S^{\prime})$.
+But since $r$ is a rotation, it’s a bijection, so $r^{-1}(S^{\prime})$ is just another set of size $|S|$.
+[Wrong] correct to wrong: The initial thought process was on the right track, but the self-correction led to incorrect reasoning steps.
+Wait, but perhaps the number is 105.
+[Wrong] Wait, but 105 is $2\cdot 10^{2}$, which is 105.
+Wait, perhaps the number is 105.
+[Wrong] wrong to wrong: Repeated incorrect self-correction, circling back to the same incorrect value.
+Case 2: Reattempt a previously tried approach.
+Wait, but in this case, it’s more specific: the set must be entirely disjoint from some rotated copy of itself.
+Hmm.
+Wait, maybe it’s useful to think in terms of necklace structures.
+A necklace is an equivalence class under rotation, but I’m not sure.
+[The method “necklace counting” was tried before and failed.] This highlights an instance where the model redundantly reapplied a previously attempted strategy it had already recognized as ineffective for the current problem.
+Case 3: Overthinking trivial details.
+Wait, but actually, for the identity rotation, $r(S)=S$, so $S\cap r(S)=S$, which is not empty unless $S$ is empty.
+So, for the identity rotation, the condition $S\cap r(S)=\emptyset$ is only satisfied when $S$ is empty.
+So, for non-empty $S$, the identity rotation cannot satisfy the condition.
+Therefore, for non-empty $S$, we need some non-identity rotation $r$ where $S\cap r(S)=\emptyset$.
+Wait, actually, the problem says ‘can be rotated’, so it includes the identity rotation.
+But as we saw, for non-empty $S$, the identity rotation only works if $S$ is empty, which is trivial.
+So, perhaps the problem is really about non-empty $S$ where some non-identity rotation makes $S\cap r(S)=\emptyset$.” This lengthy deliberation focuses on the trivial case of the identity rotation and its implications.
+This demonstrates a scenario where the model’s reflection-driven over-complication of a simple problem result in unnecessarily verbose reasoning.
+Table 5: Examples of redundant reflection behaviors extracted from the response generated by the DeepSeek-R1-Distill-Qwen-7B model on a specific problem in AIME24 dataset.
+The original problem is “Each vertex of a regular octagon is independently colored either red or blue with equal probability.
+The probability that the octagon can then be rotated so that all of the blue vertices end up at positions where there were originally red vertices is $\tfrac{m}{n}$, where $m$ and $n$ are relatively prime positive integers.
+What is $m+n$?
+Appendix D Relationship Between Certainty and Correctness
+Figure 4: Certainty distributions of probes with correct and incorrect results on AIME24, AMC23, and MATH500.
+Our further analysis correlation between reasoning certainty and correctness.
+Specifically, we collect reasoning trajectories generated by DeepSeek-R1-Distill-Qwen-7B on three benchmarks: AIME24, AMC23, and MATH500.
+For each trajectory, we insert fixed-interval probing points every 128 tokens to examine (1) the correctness of the probed content, and (2) the certainty score computed by our method.
+As shown in Figure 4, reasoning steps with high certainty (e.g., >0.9>0.9) are significantly more likely to be correct, indicating that the model’s self-estimated certainty serves as a reliable indicator of reasoning quality.